Grade 11Physical Chemistry

A beam of electron accelerated with 4.64v is passed through a tube containing mercury vapour. As a result of absorption electronic changes occured. If full energynof singlenelectron was covered into light what was wavenumber of light

Abhiram
7 Years agoGrade 11

1 Answer

Sahil Yadav
6 Years ago
Energy absorbed=4.64eV=4.64×1.6×1019J=7.424×10nJ
λhc=energy
 wave number =λ1=hcEnergyobserved=6.626×1034×3×1087.424×1019m1=3.75×104 cm1